[Ncep.nhc.nco_contacts] HSU Shift Log -- Technical Edition -- Wed Aug 17, 2016

HSU Shift Log nhc.tsbadmin at noaa.gov
Wed Aug 17 20:33:26 UTC 2016


HSU Shift Log for Wed Aug 17 2016 16:33:20 GMT-0400 (EDT)

Shift: H

Forecaster: SR Stewart

Technical Issues Summary:

ATCF issues: None.

N-AWIPS/AWIPS2 issues: None.

Other Techincal issues: Hissing sound/static remains on Hotline; was unable  
to hear WPC at times during Med Range Conf. call due to background noise.
-------------- next part --------------
An HTML attachment was scrubbed...
URL: https://www.lstsrv.ncep.noaa.gov/pipermail/ncep.nhc.nco_contacts/attachments/20160817/2f1b8216/attachment.html 


More information about the Ncep.nhc.nco_contacts mailing list